Question 1:1 Brad Vaughn Corporation sells two types of computers; one is designed for audio applications and the other for video applications Vaughn incurs $254720 in bed costs Per-unit data on the two products is presented blow Unit data Selling price Variable costs Contribution margin Sales mix Audio computer $1450 1080 $370 80% Video computer $1710 1200 $510 20% What ika be the total contribution margin at the break-even point? $147260 O $202980 $254720 $350240
Answer #1

Answer:

Option 3rd is correct. i.e $254,720

At BEP ,Fixed cost = Total Contribution

Weighted average contribution margin = (370*80%+510*20%) = 398 per units
Break even point = $254,720/398 = 640 Units
Total contribution margin at break even point is = 640*398 = $254,720